# Exact output tracking for the one-dimensional heat equation and applications to the interpolation problem in Gevrey classes of order 2

### v1

@@ Acknowledgements
The authors are deeply indebted to Oliver Glass for many valuable discussions and insightful suggestions. Lucas Davron thanks Sylvain Ervedoza for fruitful discussions and having suggested to use the Fourier-Laplace transform. The authors also thank Swann Marx for reading a preliminary version of this paper. Pierre Lissy was supported by the french ‘‘Agence Nationale de la Recherche’’ under the grant ANR-22-CPJ1-0027-01.
2 Exact tracking in infinite time horizon
To show Proposition 6 we rely on the theory of boundary values of holomorphic functions, which we recall below.
@@ 2.1 Boundary values of holomorphic functions
Let $p\in[1,+\infty[$. A function $F\in\operatorname{Hol}(\mathbb{C}_{+})$ is said to belong to $H^{p}(\mathbb{C}_{+})$ if
$\|F\|_{H^{p}(\mathbb{C}_{+})}^{p}:=\sup_{x>0}\int_{\mathbb{R}}|F(x+i\xi)|^{p}d\xi<\infty.$
A function $F\in\operatorname{Hol}(\mathbb{C}_{+})$ is said to belong to $H^{\infty}(\mathbb{C}_{+})$ if
$\|F\|_{H^{\infty}(\mathbb{C}_{+})}:=\sup_{z\in\mathbb{C}_{+}}|F(z)|<\infty.$
The above definition naturally yields a Banach space for any $p\in[1,+\infty]$. The following result is due to Paley and Wiener. In particular, it implies that $H^{2}(\mathbb{C}_{+})$ is also a Hilbert space.
@@
$\mathcal{L}:L^{2}(0,\infty)\rightarrow H^{2}(\mathbb{C}_{+}),\quad\mathcal{L}f(s)=\int_{0}^{\infty}e^{-st}f(t)dt.$
@@
The theory of boundary values for Hardy functions begins with Fatou’s Theorem. It is usually proved first for functions holomorphic on the unit disc [rudin, Theorem 17.10] and then exported to other geometries by conformal mapping.
Theorem 13.
Let $F\in H^{2}(\mathbb{C}_{+})$ and $0<\alpha<1$. Then for almost every $i\xi\in i\mathbb{R}$, the limit
$\lim_{\Gamma_{\alpha}(i\xi)\ni z\rightarrow i\xi}F(z),$
exists. The induced function $\xi\mapsto F(i\xi)$ is $L^{2}(\mathbb{R})$.
@@
Such limits are referred to as non-tangential, or sometimes angular. The next result is often called Lusin-Privalov’s uniqueness theorem.
@@
[koosis, III.D] Let $F\in\operatorname{Hol}(\mathbb{C}_{+})$, and assume that for a subset $E\subset i\mathbb{R}$ of positive measure, there holds
$\exists 0<\alpha<1,\quad\forall i\xi\in E,\quad\lim_{\Gamma_{\alpha}(i\xi)\ni z\rightarrow i\xi}F(z)=0.$
Then $F$ is identically zero on $\mathbb{C}_{+}$.
@@
[rudin, Theorem 19.3] Let $f\in L^{2}(-A,A)$ for some $A>0$. Then $\mathcal{F}f\in\operatorname{Hol}(\mathbb{C})\cap L^{2}(\mathbb{R})$ satisfies
@@
Conversely, if $\phi\in\operatorname{Hol}(\mathbb{C})\cap L^{2}(\mathbb{R})$ satisfies the above estimate, then $\mathcal{F}\phi$ has support contained in [A,A][-A,A].
@@
    There exists a function $F\in H^{2}(\mathbb{C}_{+})$ such that
@@
Combining Theorem 16 and Lusin-Privalov’s uniqueness theorem, we obtain the following characterization of the holomorphic functions belonging to $H^{2}(\mathbb{C}_{+})$, from their boundary value.
@@
Let $F\in\operatorname{Hol}(\mathbb{C}_{+})$ have boundary value $f\in L^{1}_{\operatorname{loc}}(\mathbb{R})$ in the following sense: there is $0<\alpha<1$ such that for almost every $\xi\in\mathbb{R}$
$\lim_{\Gamma_{\alpha}(i\xi)\ni z\rightarrow i\xi}F(z)=f(\xi).$
Then $F\in H^{2}(\mathbb{C}_{+})$ if and only if $f$ satisfies both
@@ 2.2 Spectrum of entire functions
We are now in a position to prove Proposition 6. We will need the following definition.
@@
The order $\rho$ of an entire non-constant function $f$ is defined as
@@
If $\rho<\infty$, the type $\tau$ of $f$ is defined by
$\tau=\limsup_{r\to\infty}\frac{\log M_{f}(r)}{r^{\rho}}.$
In particular, if both $\rho$ and $\tau$ are finite, we have
$\forall\epsilon>0,\quad\exists C>0,\quad\forall z\in\mathbb{C},\quad|f(z)|\leq e^{(\tau+\epsilon)|z|^{\rho}},$
and $\tau$ is the best such constant. It is customary to say that $f$ has exponential type $\leq\tau$ when the above inequality holds with $\rho=1$.
@@
Recall that Lemma 19 is motivated by the formal computation
@@
which uses $\operatorname{supp}\mathcal{F}\phi\subset\{0\}$. However, as already pointed out in the introduction, for $\phi$ entire of order <1<1, the restriction of $\phi$ to the real axis might not be a tempered distribution. We can give a meaning to $\mathcal{F}\phi$ using Gelfand–Silov’s generalized function space $(S_{1}^{1})^{\prime}$ or the theory of hyperfunctions, see [gelfand_silov, gs2, roumieu, smirnov, gl_eng, kaneko]. However, in this setting it is not clear how to define the convolution $\mathcal{F}\phi*\mathcal{F}f$, hence we are not able to directly use the above formal computation.
@@
Now consider $\phi_{\epsilon}:=(\mathcal{F}\varrho_{\epsilon})\phi\in\operatorname{Hol}(\mathbb{C})$. On the one hand, from (21) and being $\phi$ of order <1<1 we get
@@
for some constants $C,R>0$ and $0<\rho<1$, not depending on $\xi\in\mathbb{R}$. Taking $1<s<1/\rho$ ensures that the above function is square-integrable over the real line hence $\phi_{\epsilon}\in L^{2}(\mathbb{R})$. On the other hand, because $\phi$ has order $\rho<1$ and (22), it follows that $\phi_{\epsilon}$ is of exponential type $\leq\epsilon$. Using again the Paley–Wiener theorem we obtain that $\mathcal{F}\phi_{\epsilon}$ is supported in $\epsilon,\epsilon]$, which implies that
@@
The given proof constructs a family $(\sigma_{\epsilon})_{\epsilon>0}$ of multipliers for the holomorphic function $\phi$, with the following properties
$\sigma_{\epsilon}\in\operatorname{Hol}(\mathbb{C})\cap L^{2}(\mathbb{R}),\quad\sigma_{\epsilon}\phi\in L^{2}(\mathbb{R}),\quad\sigma_{\epsilon}\phi\xrightarrow[\epsilon\rightarrow 0^{+}]{\mathcal{D}^{\prime}(\mathbb{R})}\phi,\quad\operatorname{supp}\mathcal{F}\sigma_{\epsilon}\subset[-\epsilon,\epsilon].$ (23)
An interesting question is whether this construction can be adapted to the case where one only assumes that $\phi$ is of exponential type 0 on $\mathbb{C}$. If $\phi$ is non-constant, a necessary condition for (23) is that the logarithmic integral converges, i.e.
@@
see [K2, §X.A]. Conversely, from the Beurling–Malliavin Theorem [K2, p. 397], if $\phi$ has finite exponential type then (24) implies that for every $\epsilon>0$, there exists some $\sigma_{\epsilon}$ satisfying every the properties in (23), except possibly the convergence of $\sigma_{\epsilon}\phi$ to $\phi$. We do not know if the condition (24), together with the assumption of $\phi$ being of exponential type 0, implies the existence of $\sigma_{\epsilon}$ such that (23) holds, for every $\epsilon>0$.
@@ 2.4 Proof of the main results
In what follows every the signals defined on $(0,\infty)$ are extended by 0 on $(-\infty,0)$. Assume first that $y\in\mathcal{Y}(0,\infty)$, let us show that $y$ is Laplace transformable and its Laplace transform satisfies (12). A formal proof is done by taking the Laplace transform in time in (1) (without checking whether it is legitimate or not), solving the resulting second order differential equation in space, using the boundary conditions given by the second and third lines of (1) and then considering $x=0$.
Let us now give a rigorous proof. We invoke the representation formula
$z(t,x)=2\int_{0}^{t}\theta(x-1,t-\sigma)u(\sigma)d\sigma,$ (32)
where
$\theta(t,x)=\sum_{m=-\infty}^{+\infty}K(t,x+2m),\quad K(t,x)=\frac{e^{-\frac{x^{2}}{4t}}}{\sqrt{4\pi t}}.$
Such a formula can be found e.g. in [cannon, Theorem 6.4.1]. Evaluating (32) at $x=0$, we obtain
$y(t)=z(t,0)=\int_{0}^{t}k(t-\sigma)u(\sigma)d\sigma,$ (33)
where we have introduced the kernel
$k(t)=\sum_{m=-\infty}^{+\infty}K(t,2m-1)=\frac{2}{\sqrt{\pi t}}\sum_{m=-\infty}^{+\infty}e^{-\frac{(m+1/2)^{2}}{t}},\,t>0,$
extended by 0 on $(-\infty,0]$. Crude estimates show that
$\exists C>0,\quad\forall t>0,\quad|k(t)|\leq C\sqrt{t},$
so that
$|y(t)|\lesssim\|u\|_{L^{2}(0,\infty)}t,$
hence $y$ is Laplace transformable, with abscissa of absolute convergence $\sigma_{y}\leq 0$. Since $k$ and $u$ are supported in $0,\infty)$, we recognize this as the usual convolution product, and taking the Laplace transform, we have
$\hat{y}(s)=\hat{k}(s)\hat{u}(s).$
An explicit computation gives that
@@
We deduce that the transfer function of (1) is given by
@@
Note that $1/\phi$ is bounded on $\mathbb{C}_{+}$, so that $\hat{\dot{y}}\in H^{2}(\mathbb{C}_{+})$ and $\dot{y}\in L^{2}(0,\infty)$. In particular, the Laplace transform $\hat{\dot{y}}$ has boundary value $\hat{\dot{y}}(i\xi)=\mathcal{F}\dot{y}(\xi)$. Now, coming back to
$\phi\hat{\dot{y}}=\hat{u}\in H^{2}(\mathbb{C}_{+}),$
@@
where $z$ is the solution of (1) with control $u$. In particular, we have $\partial_{t}^{k}z(T,0)=\varphi^{(k)}(T^{-})$ for every $k\in\mathbb{N}$. We then extend the control $u$ by 0 on $(T,\infty)$, which extends $z$ for times $t>T$. Consider then

### v2

@@ Acknowledgements
The authors are deeply indebted to Oliver Glass for many valuable discussions and insightful suggestions. Lucas Davron thanks Sylvain Ervedoza for fruitful discussions and having suggested to use the Fourier-Laplace transform. The authors also thank Swann Marx for reading a preliminary version of this paper. Pierre Lissy was supported by the french “Agence Nationale de la Recherche” under the grant ANR-22-CPJ2-0138-01.
2 Exact tracking
To show Proposition 6 we will rely on the theory of boundary values of holomorphic functions, which we recall below.
@@ 2.1 Boundary values of holomorphic functions
Let $p\in[1,+\infty[$. A function $F\in\mathcal{H}(\mathbb{C}_{+})$ is said to belong to $\mathcal{H}^{p}(\mathbb{C}_{+})$ if
$\|F\|_{\mathcal{H}^{p}(\mathbb{C}_{+})}^{p}:=\sup_{x>0}\int_{\mathbb{R}}|F(x+i\xi)|^{p}d\xi<\infty.$
A function $F\in\mathcal{H}(\mathbb{C}_{+})$ is said to belong to $\mathcal{H}^{\infty}(\mathbb{C}_{+})$ if
$\|F\|_{\mathcal{H}^{\infty}(\mathbb{C}_{+})}:=\sup_{z\in\mathbb{C}_{+}}|F(z)|<\infty.$
The above definition naturally yields a Banach space for any $p\in[1,+\infty]$. The following result is due to Paley and Wiener. In particular, it implies that $\mathcal{H}^{2}(\mathbb{C}_{+})$ is also a Hilbert space.
@@
$\mathcal{L}:L^{2}(0,\infty)\rightarrow\mathcal{H}^{2}(\mathbb{C}_{+}),\quad\mathcal{L}f(s)=\int_{0}^{\infty}e^{-st}f(t)dt.$
@@
Definition 13.
Let $F:\mathbb{C}_{+}\rightarrow\mathbb{C}$ holomorphic and $f:\mathbb{R}\rightarrow\mathbb{C}$ measurable. We say that $F$ has the non tangential (or angular) boundary value $f$ if there exists a constant $0<\alpha<1$ and a subset $E\subset\mathbb{R}$ of positive measure such that
$\forall\xi\in E,\quad F(z)\xrightarrow[z\rightarrow i\xi]{z\in\Gamma_{\alpha}(i\xi)}f(\xi).$
@@
The set $E$ is authorized to merely be of positive measure from Luzin-Privalov’s uniqueness theorem, which is stated below. This result, as well as Fatou’s Theorem 15 below, is usually proved first for functions holomorphic on the unit disc and then exported to other geometries by conformal mapping.
@@
[koosis, III.D] Let $F\in\mathcal{H}(\mathbb{C}_{+})$, and assume that $F$ has angular boundary value zero. Then $F$ is identically zero on $\mathbb{C}_{+}$.
We will frequently abuse notations and write $F(i\xi)$ for $f(\xi)$. Note also that the function $f$ may be defined up to a zero measure set.
The next result is an existence theorem for boundary values, due to Fatou.
Theorem 15.
[rudin, Theorem 17.10] Let $F\in\mathcal{H}^{2}(\mathbb{C}_{+})$ and $0<\alpha<1$. Then for almost every $\xi\in\mathbb{R}$, the limit
$\lim_{\Gamma_{\alpha}(i\xi)\ni z\rightarrow i\xi}F(z)=:F(i\xi),$
exists. The induced function $\xi\mapsto F(i\xi)$ is $L^{2}(\mathbb{R})$.
@@
[rudin, Theorem 19.3] Let $f\in L^{2}(-A,A)$ for some $A>0$. Then $\mathcal{F}f\in\mathcal{H}(\mathbb{C})\cap L^{2}(\mathbb{R})$ satisfies
@@
Conversely, if $\phi\in\mathcal{H}(\mathbb{C})\cap L^{2}(\mathbb{R})$ satisfies the above estimate, then $\mathcal{F}\phi$ has support contained in [A,A][-A,A].
@@
    There exists a function $F\in\mathcal{H}^{2}(\mathbb{C}_{+})$ such that
@@
Combining Theorem 17 and Lusin-Privalov’s uniqueness theorem, we obtain the following characterization of the holomorphic functions belonging to $\mathcal{H}^{2}(\mathbb{C}_{+})$, from their boundary value.
@@
Let $F\in\mathcal{H}(\mathbb{C}_{+})$ and $f:\mathbb{R}\rightarrow\mathbb{C}$ measurable, possibly defined up to a zero measure set. Assume that $F$ has angular boundary value $f$. Then $F\in\mathcal{H}^{2}(\mathbb{C}_{+})$ if and only if $f$ satisfies both
@@ 2.2 Spectrum of entire functions
We are now in position to prove Proposition 6. We will need the following definition.
@@
[boas, Chapter 2] The order $\rho$ of an entire non-constant function $f$ is defined as
@@
If $\rho<\infty$, the type $A$ of $f$ is defined by
$A=\limsup_{r\to\infty}\frac{\log M_{f}(r)}{r^{\rho}}.$
In particular, if both $\rho$ and $A$ are finite, we have
$\forall\epsilon>0,\quad\exists C>0,\quad\forall z\in\mathbb{C},\quad|f(z)|\leq e^{(A+\epsilon)|z|^{\rho}},$
and $A$ is the best such constant.
@@
Lemma 20 is formally shown with the computations
@@
which uses the formal assertion $\operatorname{supp}\mathcal{F}\phi\subset\{0\}$. We can give a meaning to $\mathcal{F}\phi$ using Gelfand–Silov’s generalized function space $(S_{1}^{1})^{\prime}$ or the theory of hyperfunctions, see [gs2, kaneko]. However, in this setting it is not clear for us how to define the convolution $\mathcal{F}\phi*\mathcal{F}f$, which prevents us to use the associated Paley-Wiener theorem [roumieu, Théorème 1, §III.1, p. 98], [gl_eng, Chapter 2, p. 259].
@@
Now consider $\phi_{\epsilon}:=(\mathcal{F}\varrho_{\epsilon})\phi\in\mathcal{H}(\mathbb{C})$. On the one hand, from (21) and being $\phi$ of order <1<1 we get
@@
for some constants $C,R>0$ and $0<\rho<1$, not depending on $\xi\in\mathbb{R}$. Taking $1<s<1/\rho$ ensures that the above function is square-integrable over the real line, whence $\phi_{\epsilon}\in L^{2}(\mathbb{R})$. On the other hand, because $\phi$ has order $\rho<1$ and (22), it follows that $\phi_{\epsilon}$ is of exponential type $\leq\epsilon$. Using again the Paley–Wiener theorem we obtain that $\mathcal{F}\phi_{\epsilon}$ is supported in $\epsilon,\epsilon]$, which implies that
@@
The given proof constructs a family $(\sigma_{\epsilon})_{\epsilon>0}$ of multipliers for the holomorphic function $\phi$, such that for all $\epsilon>0$ we have
$\sigma_{\epsilon}\in\mathcal{H}(\mathbb{C})\cap L^{2}(\mathbb{R}),\quad\sigma_{\epsilon}\phi\in L^{2}(\mathbb{R}),\quad\operatorname{supp}\mathcal{F}(\sigma_{\epsilon}\phi)\subset[-\epsilon,\epsilon],$ (23)
together with
$\forall f\in L^{2}(\mathbb{R}),\quad\phi f\in L^{2}(\mathbb{R})\Longrightarrow\sigma_{\epsilon}\phi f\xrightarrow[\epsilon\rightarrow 0^{+}]{L^{2}(\mathbb{R})}\phi f.$ (24)
This allows to give a definition of $\operatorname{supp}\mathcal{F}\phi\subset\{0\}$, which is moreover compatible with the standard distributional calculus. It should be noticed that if $\phi$ has no growth restriction on the real axis then its Fourier transform, as a generalized function, may not be supported on the real line. For instance, in [gl_eng, Chapter 2] it is shown that the Fourier transform of $e^{x}$ can be identified as the analytic functional which is the evaluation at $z=i$. Moreover, a necessary condition for the existence of $\sigma_{\epsilon}$ such that (23) holds is that the logarithmic integral converges, i.e.
@@
see [K2, §X.A]. Conversely, if $\phi$ has finite exponential type and satisfies (25), then from the Beurling–Malliavin Theorem [K2, p. 397] there exists some $\sigma_{\epsilon}$ satisfying (23), for all $\epsilon>0$. We do not know if one may chose $\sigma_{\epsilon}$ so as (24) is also satisfied.
@@ 2.4 Proof of the main results
For the system 1 we consider transposition solutions [Coron, §2.3].
Definition 22.
Let $u\in L^{2}(0,\infty)$, a solution of (1) is $z\in C([0,\infty);L^{2}(0,1))$ such that for all $0<T<\infty$ and all $\varphi^{T}\in L^{2}(0,1)$, denoting $\varphi(t,x)$ the solution of
$\left\{\begin{array}[]{rclcc}-\varphi_{t}(t,x)&=&\varphi_{xx}(t,x),&0<t<T,&0<x<1,\\ \varphi_{x}(t,1)&=&0,\\ \varphi_{x}(t,0)&=&0,\\ \varphi(T,x)&=&\varphi^{T}(x),\end{array}\right$.
we have
$\int_{0}^{1}z(T,x)\varphi^{T}(x)dx=\int_{0}^{T}\varphi(t,1)u(t)dt.$
It is easy to see that (1) is then well-posed in the sense that for all $u\in L^{2}(0,\infty)$, there exists a unique solution $z$ to (1), and that
$\forall 0<T<\infty,\quad\exists C>0,\quad\forall u\in L^{2}(0,\infty),\quad\|z\|_{C([0,T];L^{2}(0,1))}\leq C\|u\|_{L^{2}(0,T)}.$
Moreover, introducing the spectral decomposition of the Neumann Laplacian
$\lambda_{j}=(j\pi)^{2},\quad e_{j}(x)=\left\{\begin{array}[]{cc}1&j=0,\\ \sqrt{2}\cos(j\pi x)&j\geq 1,\end{array}\right.,\quad(j\in\mathbb{N})$
one easily verifies that $z$ assumes the representation
$z(t)=\int_{0}^{t}\sum_{j=0}^{\infty}e^{-\lambda_{j}(t-\sigma)}e_{j}(1)u(\sigma)e_{j}d\sigma,$ (33)
where for fixed $t$ there holds
$\int_{0}^{t}\sum_{j=0}^{\infty}\left\|e^{-\lambda_{j}(t-\sigma)}e_{j}(1)u(\sigma)e_{j}\right\|_{C[0,1]}d\sigma<\infty.$
Thus, the function $z$ belongs to $C([0,\infty)\times[0,1])$, and the output $y(t)$ makes sense as the pointwise evaluation of $z(t,x)$ at $x=0$. In fact, using the hypoellipticity of the heat operator $\partial_{t}-\partial_{xx}$, one can further show that $y\in C^{\infty}_{(0)}[0,\infty)$.
@@
In what follows all the signals defined on $(0,\infty)$ are extended by 0 on $(-\infty,0)$. Assume first that $y\in\mathcal{Y}(0,\infty)$, let us show that $y$ is Laplace transformable and its Laplace transform satisfies (12). A formal proof is done by taking the Laplace transform in time in (1) (without checking whether it is legitimate or not), solving the resulting second order differential equation in space, using the boundary conditions given by the second and third lines of (1) and finally evaluating at $x=0$.
The above method will be applied in §3.1.1 to another system. For (1) it will be more convenient to work with the representation formula (33). The representation formula (33) may be evaluated at $x=0$, yielding
$y(t)=z(t,0)=\int_{0}^{t}k(t-\sigma)u(\sigma)d\sigma,\quad k(t):=1+2\sum_{j=1}^{\infty}e^{-(j\pi)^{2}t}(-1)^{j},$
where the defined kernel $k$ is extended by 0 for $t<0$. The function $k$ may be rewritten using the Poisson summation formula as
$k(t)=\frac{1}{\sqrt{\pi t}}\sum_{m=-\infty}^{+\infty}e^{-\frac{(m+1/2)^{2}}{t}}.$
Combining the two representations of $k$, we arrive to $k\in C^{\infty}_{(0)}[0,\infty)\cap L^{\infty}(0,\infty)$. The last fact can be used as an alternative proof of $y\in C_{(0)}^{\infty}[0,\infty)$. The signals $k,u,y$ are therefore all Laplace transformable, with abscissa of absolute convergence no greater than 0. The convolution rule further applies:
$\hat{y}(s)=\hat{k}(s)\hat{u}(s),\quad s\in\mathbb{C}_{+},$
where an explicit computation gives that
@@
In the above computation, the interversion series-integral is justified by considering first $s>0$, which yields only positive quantities in the summand. We deduce that the transfer function of (1) is given by
@@
Note that $1/\phi$ is bounded on $\mathbb{C}_{+}$, so that $\hat{\dot{y}}\in\mathcal{H}^{2}(\mathbb{C}_{+})$ and $\dot{y}\in L^{2}(0,\infty)$. In particular, the Laplace transform $\hat{\dot{y}}$ has the non tangential boundary value $\hat{\dot{y}}(i\xi)=\mathcal{F}\dot{y}(\xi)$. Now, coming back to
$\phi\hat{\dot{y}}=\hat{u}\in\mathcal{H}^{2}(\mathbb{C}_{+}),$
@@
where $z$ is the solution of (1) with control $u$. In particular, we have $\partial_{t}^{k}z(T,0)=\varphi^{(k)}(T^{-})$ for every $k\in\mathbb{N}$. We then extend the control $u$ by 0 on $(T,\infty)$, which extends $z$ for times $t>T$ as a $C^{\infty}([0,\infty)\times(-1,1))$. Consider then
@@
Remark 23.
For an initial data $z^{0}\in L^{2}(0,1)$, the solution $z$ of
$\left\{\begin{array}[]{rclcc}z_{t}(t,x)&=&z_{xx}(t,x),&0<x<1,&t>0,\\ z_{x}(t,1)&=&0,\\ z_{x}(t,0)&=&0,\\ z(0,x)&=&z^{0}(x),\end{array}\right$. (36)
is of class $C^{\infty}((0,\infty)\times(-1,1))$ and the output $y(t)=z(t,0)$ is in $L^{p}(0,T)$ for all $1\leq p<4$ and $0<T<\infty$. The value of $p$ cannot be increased to $p>4$. This however has no impact on the proof of Theorem 2, as therein we only consider such outputs for initial data in the reachable space, which are seen to be exactly shifted outputs of the system (1) (with a control and zero initial data). In other words, we do not need the existence of the output $y$ for (36).